OpenStudy (anonymous):

factor completely. x^2+7xy-30y^2

OpenStudy (anonymous):

What can you multiply and get -30 and add together to get 7

OpenStudy (anonymous):

1*-30=-30 10-3=7 10*-3=-30 write 7=(10-3) and then make factors or 7xy=10xy-3xy and make factors.